A radio station's channel, such as 100.7 FM or 92.3 FM, is actually its frequency in megahertz... a. A radio station's channel, such as 100.7 FM or 92.3 FM, is actually its frequency in megahertz (Mhz), where 1 MH is 10^6 H. Calculate the broadcast wavelength of the radio station 88.30 FM
Answer #1

f = 88.30 FM

= 88.30 MHz

= 88.30*10^6 Hz

= 8.83*10^7 Hz

= 8.83*10^7 s-1

use:

lambda = c/f

=(3.0*10^8 m/s)/(8.83*10^7 s-1)

= 3.40 m

Answer: 3.40 m